# CCMNet: Leveraging Calibrated Color Correction Matrices for Cross-Camera Color Constancy

### v1

@@ 4.2 Results
Figure 5: Visual comparison of the results from C5 [6] with different additional image sets (second, third column) and CCMNet (fourth column). While C5 relies on additional images, CCMNet is optimized for fixed CFE guidance, ensuring consistent performance.
@@
Gehler-Shi [57] Mean Med. Tri. B.25% W.25% Size(MB)
2nd-order Gray-Edge [61] 5.13 4.44 4.62 2.11 9.26 -
Shades-of-Gray [23] 4.93 4.01 4.23 1.14 10.20 -
PCA-based B/W Colors [19] 3.52 2.14 2.47 0.50 8.74 -
ASM [8] 3.80 2.40 2.70 - - -
Woo et al. [62] 4.30 2.86 3.31 0.71 10.14 -
Grayness Index [56] 3.07 1.87 2.16 0.43 7.62 -
Cross-dataset CC [43] 2.87 2.21 - - - -
Quasi-Unsupervised CC [14] 3.46 2.23 - - - 622
SIIE [3] 2.77 1.93 - 0.55 6.53 10.3
FFCC [13] 2.95 2.19 2.35 0.57 6.75 0.22
C5 ($m=7$) [6] 2.36 1.61 1.74 0.44 5.60 1.74
C5 ($m=9$) [6] 2.50 1.99 2.03 0.53 5.46 2.09
@@
Cube+ [9] Mean Med. Tri. B.25% W.25%
Gray-world [18] 3.52 2.55 2.82 0.60 7.98
1st-order Gray-Edge [61] 3.06 2.05 2.32 0.55 7.22
2nd-order Gray-Edge [61] 3.28 2.34 2.58 0.66 7.44
Shades-of-Gray [23] 3.22 2.12 2.44 0.43 7.77
Cross-dataset CC [43] 2.47 1.94 - - -
Quasi-Unsupervised CC [14] 2.69 1.76 2.00 0.49 6.45
SIIE [3] 2.14 1.44 - 0.44 5.06
FFCC [13] 2.69 1.89 2.08 0.46 6.31
DMCC [66] 2.23 1.63 1.78 0.49 4.95
C5 ($m=7$) [6] 1.87 1.27 1.40 0.41 4.36
C5 ($m=9$) [6] 1.92 1.32 1.46 0.44 4.44
@@
NUS-8 [19] Mean Med. Tri. B.25% W.25%
Gray-world [18] 4.59 3.46 3.81 1.16 9.85
Shades-of-Gray [23] 3.67 2.94 3.03 0.98 7.75
Local Surface Reflectance [28] 3.45 2.51 2.70 0.98 7.32
PCA-based B/W Colors [19] 2.93 2.33 2.42 0.78 6.13
Grayness Index [56] 2.91 1.97 2.13 0.56 6.67
Cross-dataset CC [43] 3.08 2.24 - - -
Quasi-Unsupervised CC [14] 3.00 2.25 - - -
FFCC [13] 2.87 2.14 2.30 0.71 6.23
C5 ($m=7$) [6] 2.68 2.00 2.14 0.66 5.90
C5 ($m=9$) [6] 2.54 1.90 2.02 0.61 5.61
@@
NUS-8 (CS) [19] Mean Med. Tri. B.25% W.25%
DMCC (CS) [66] 2.80 2.12 2.25 0.74 5.88
SIIE (CS) [3] 2.05 1.50 - 0.52 4.48
C5 ($m=9$, CS) [6] 1.77 1.37 1.46 0.48 3.75
@@
We also report results on the cross-sensor (CS) validation setup [3]. In this protocol, the network is trained on data from seven cameras in the NUS-8 dataset, excluding one as the test camera. This process is repeated for each of the eight cameras, and the results are averaged.
@@ 4.3 Generalization with SIIE Backbone
We further explore using CFE with SIIE [3]. SIIE learns a 3$\times$3 matrix by processing the raw image $uv$-histogram to map raw colors to a color working space. In this experiment, we replace C5 with SIIE as our backbone. Specifically, we input our CFE-concatenated $uv$-histograms, augmented with the imaginary camera transformation, into the SIIE backbone. As shown in Table 2, the best performance (MAE) is achieved when both CFE and augmentation are applied, confirming that CCMNet generalizes to different backbones utilizing $uv$-histograms.
@@
Table 2: Generalization with the SIIE [3] backbone. Reported results show the mean angular error.
@@ 4.4 Ablation Studies
Table 3 presents the performance of the backbone $f$ and CCMNet trained under three setups: without augmentation (w/o aug), with augmentation at $\alpha=1$, and with augmentation for $0\leq\alpha\leq 1$. Architecturally, the backbone $f$ mirrors the $m=1$ structure from C5 [6], excluding additional images and encoders. Setting $\alpha=1$ in the augmentation process replicates the camera-mapping strategy used in C5.
@@ Appendix B Details of the CFE Encoding Process
As mentioned in the main paper, the sampled XYZ locus was transformed into the camera’s native raw RGB space by interpolating between CM1 and CM2. This was further converted into a histogram with 64 bins, within the $uv$ range of [-0.5, 1.5]. The resulting 64$\times$64$\times$1 histogram was processed by the CFE encoder, which outputs an 8-dimensional embedding vector. The CFE encoder consists of four DoubleConvBlocks followed by a projection head. Each DoubleConvBlock processes the input by applying two convolutional layers, each with a kernel size of $3\times 3$, a stride of 1, and a padding of 1, followed by a LeakyReLU activation. This is then followed by a $2\times 2$ max-pooling layer and batch normalization. The projection head flattens the feature map and maps it to an 8-dimensional embedding vector using an MLP with two hidden layers.
@@ Appendix C Camera-to-Camera Mapping
In Sec. 3.4 of the main paper, we introduced our imaginary camera augmentation, which assumes two versions of the same image in the camera’s native raw RGB space. To satisfy this condition, we perform a camera-to-camera mapping inspired by [6]. In this section, we provide a detailed explanation of the camera-to-camera mapping process used in our work. Specifically, in Sec. C.1, we explain the process of computing the correlated color temperature (CCT) of a light source in the target camera’s native raw RGB space. Then, in Sec. C.2, we describe how to generate a pool of white-balanced, camera-independent XYZ images using the RGB values of the light source and the corresponding CCT. In Sec. C.3, we describe the process of generating a device-specific illumination pool for random sampling. Finally, Sec. C.4 explains our camera-to-camera mapping, which presents a reference image in two different camera-native raw RGB spaces. The reference image is sampled from the XYZ image pool, while the illumination is sampled from the augmented ground-truth (GT) illumination pool of each camera. The overall process is visualized in Fig. 8.
@@
While our camera-to-camera mapping is inspired by the C5 augmentation approach [6], it differs in the following ways. First, we remove C5’s restriction that limits sampling from the illumination pool to similar scenes with matching capturing settings (e.g., ISO, exposure time) and illumination CCT. Specifically, in C5, both the sampled scene image from the CIE XYZ space and the sampled illuminant from the target camera were required to have similar capturing settings and CCT. In contrast, our approach removes this constraint, eliminating the need to rely on capturing settings and allowing for greater diversity in augmentation. Additionally, instead of sampling from a fitted cubic polynomial based on the target camera’s illuminant samples, we use a fitted cubic polynomial based on the illuminant values from the source camera’s dataset (i.e., the camera from which the reference XYZ image was taken). The sampled illuminant is then transferred to the CIE XYZ space using the inverse of the source camera’s CM, followed by a transformation of these CIE XYZ illuminant values into the native raw RGB space of the target camera.
@@ C.3 Camera-specific Illumination Pool Generation
Next, we generate an illumination pool for each camera. While it is possible to use only the GT illuminations, we adopt the augmentation method proposed in [6] to enhance generality and diversity. This method involves fitting a cubic polynomial to the GT illuminations for each camera and then introducing random shifts to augment the illuminations. For further details, please refer to the supplementary material of [6]. On the right side of Fig. 8-(B), we show a plot of the illumination pool for a specific camera (Camera A). In this plot, the red points represent the GT illumination labels extracted from the dataset, while the blue points correspond to the augmented illuminations.
@@ Appendix E Experimental Setup
As mentioned in the main paper, the backbone $f$ uses the standard U-Net-like architecture from C5 [6]. However, unlike C5, we do not use additional images from the test camera, so no extra encoders are employed. Instead, we use a single Encoder-Decoder U-Net architecture. The encoder and decoder are connected via skip connections, with each consisting of four DoubleConv layers. In the encoder, each DoubleConv layer is followed by max pooling, while in the decoder, feature upsampling and skip connections are applied before each DoubleConv layer.
The batch size was set to 16, and training was conducted over 50 epochs with an initial learning rate of $5\times 10^{-4}$. A learning rate decay of 0.5 was applied at epoch 25. The Adam optimizer [42] was used for training.
For data augmentation, camera-to-camera mapping and imaginary camera augmentation are applied exclusively using the camera subsets from the training datasets, excluding the test dataset. For instance, when evaluating the Cube+ dataset, the augmented dataset used for model training is generated from images and CCMs from the Gehler-Shi [57], NUS-8 [19], and Intel-TAU [44] datasets.

### v2

@@ 4.2 Results
Figure 5: Visual comparison of the results from C5 [afifi2021cross] with different additional image sets (second, third column) and CCMNet (fourth column). While C5 relies on additional images, CCMNet is optimized for fixed CFE guidance, ensuring consistent performance.
@@
Gehler-Shi [shi2000re] Mean Med. Tri. B.25% W.25% Size(MB)
2nd-order Gray-Edge [van2007edge] 5.13 4.44 4.62 2.11 9.26 -
Shades-of-Gray [finlayson2004shades] 4.93 4.01 4.23 1.14 10.20 -
PCA-based B/W Colors [cheng2014illuminant] 3.52 2.14 2.47 0.50 8.74 -
ASM [Akbarinia2018ColourCB] 3.80 2.40 2.70 - - -
Woo et al. [woo2017improving] 4.30 2.86 3.31 0.71 10.14 -
Grayness Index [qian2019finding] 3.07 1.87 2.16 0.43 7.62 -
Cross-dataset CC [koskinen122020cross] 2.87 2.21 - - - -
Quasi-Unsupervised CC [bianco2019quasi] 3.46 2.23 - - - 622
SIIE [afifi2019sensor] 2.77 1.93 - 0.55 6.53 10.3
FFCC [barron2017fast] 2.95 2.19 2.35 0.57 6.75 0.22
C5 ($m=7$) [afifi2021cross] 2.36 1.61 1.74 0.44 5.60 1.74
C5 ($m=9$) [afifi2021cross] 2.50 1.99 2.03 0.53 5.46 2.09
@@
Cube+ [banic2017unsupervised] Mean Med. Tri. B.25% W.25%
Gray-world [buchsbaum1980spatial] 3.52 2.55 2.82 0.60 7.98
1st-order Gray-Edge [van2007edge] 3.06 2.05 2.32 0.55 7.22
2nd-order Gray-Edge [van2007edge] 3.28 2.34 2.58 0.66 7.44
Shades-of-Gray [finlayson2004shades] 3.22 2.12 2.44 0.43 7.77
Cross-dataset CC [koskinen122020cross] 2.47 1.94 - - -
Quasi-Unsupervised CC [bianco2019quasi] 2.69 1.76 2.00 0.49 6.45
SIIE [afifi2019sensor] 2.14 1.44 - 0.44 5.06
FFCC [barron2017fast] 2.69 1.89 2.08 0.46 6.31
DMCC [Yue2023EffectiveCC] 2.23 1.63 1.78 0.49 4.95
C5 ($m=7$) [afifi2021cross] 1.87 1.27 1.40 0.41 4.36
C5 ($m=9$) [afifi2021cross] 1.92 1.32 1.46 0.44 4.44
@@
NUS-8 [cheng2014illuminant] Mean Med. Tri. B.25% W.25%
Gray-world [buchsbaum1980spatial] 4.59 3.46 3.81 1.16 9.85
Shades-of-Gray [finlayson2004shades] 3.67 2.94 3.03 0.98 7.75
Local Surface Reflectance [Gao2014EfficientCC] 3.45 2.51 2.70 0.98 7.32
PCA-based B/W Colors [cheng2014illuminant] 2.93 2.33 2.42 0.78 6.13
Grayness Index [qian2019finding] 2.91 1.97 2.13 0.56 6.67
Cross-dataset CC [koskinen122020cross] 3.08 2.24 - - -
Quasi-Unsupervised CC [bianco2019quasi] 3.00 2.25 - - -
FFCC [barron2017fast] 2.87 2.14 2.30 0.71 6.23
C5 ($m=7$) [afifi2021cross] 2.68 2.00 2.14 0.66 5.90
C5 ($m=9$) [afifi2021cross] 2.54 1.90 2.02 0.61 5.61
@@
NUS-8 (CS) [cheng2014illuminant] Mean Med. Tri. B.25% W.25%
DMCC (CS) [Yue2023EffectiveCC] 2.80 2.12 2.25 0.74 5.88
SIIE (CS) [afifi2019sensor] 2.05 1.50 - 0.52 4.48
C5 ($m=9$, CS) [afifi2021cross] 1.77 1.37 1.46 0.48 3.75
@@
We also report results on the cross-sensor (CS) validation setup [afifi2019sensor]. In this protocol, the network is trained on data from seven cameras in the NUS-8 dataset, excluding one as the test camera. This process is repeated for each of the eight cameras, and the results are averaged.
@@ 4.3 Generalization with SIIE Backbone
We further explore using CFE with SIIE [afifi2019sensor]. SIIE learns a $\times3$ matrix by processing the raw image $uv$-histogram to map raw colors to a color working space. In this experiment, we replace C5 with SIIE as our backbone. Specifically, we input our CFE-concatenated $uv$-histograms, augmented with the imaginary camera transformation, into the SIIE backbone. As shown in Table 2, the best performance (MAE) is achieved when both CFE and augmentation are applied, confirming that CCMNet generalizes to different backbones utilizing $uv$-histograms.
@@
Table 2: Generalization with the SIIE [afifi2019sensor] backbone. Reported results show the mean angular error.
@@ 4.4 Ablation Studies
Table 3 presents the performance of the backbone $f$ and CCMNet trained under three setups: without augmentation (w/o aug), with augmentation at $\alpha=1$, and with augmentation for $0\leq\alpha\leq 1$. Architecturally, the backbone $f$ mirrors the $m=1$ structure from C5 [afifi2021cross], excluding additional images and encoders. Setting $\alpha=1$ in the augmentation process replicates the camera-mapping strategy used in C5.
@@ Appendix B Details of the CFE Encoding Process
As mentioned in the main paper, the sampled XYZ locus was transformed into the camera’s native raw RGB space by interpolating between CM1 and CM2. This was further converted into a histogram with 64 bins, within the $uv$ range of [-0.5, 1.5]. The resulting $\times64$$\times$1 histogram was processed by the CFE encoder, which outputs an 8-dimensional embedding vector. The CFE encoder consists of four DoubleConvBlocks followed by a projection head. Each DoubleConvBlock processes the input by applying two convolutional layers, each with a kernel size of $3\times 3$, a stride of 1, and a padding of 1, followed by a LeakyReLU activation. This is then followed by a $2\times 2$ max-pooling layer and batch normalization. The projection head flattens the feature map and maps it to an 8-dimensional embedding vector using an MLP with two hidden layers.
@@ Appendix C Camera-to-Camera Mapping
In Sec. 3.4 of the main paper, we introduced our imaginary camera augmentation, which assumes two versions of the same image in the camera’s native raw RGB space. To satisfy this condition, we perform a camera-to-camera mapping inspired by [afifi2021cross]. In this section, we provide a detailed explanation of the camera-to-camera mapping process used in our work. Specifically, in Sec. C.1, we explain the process of computing the correlated color temperature (CCT) of a light source in the target camera’s native raw RGB space. Then, in Sec. C.2, we describe how to generate a pool of white-balanced, camera-independent XYZ images using the RGB values of the light source and the corresponding CCT. In Sec. C.3, we describe the process of generating a device-specific illumination pool for random sampling. Finally, Sec. C.4 explains our camera-to-camera mapping, which presents a reference image in two different camera-native raw RGB spaces. The reference image is sampled from the XYZ image pool, while the illumination is sampled from the augmented ground-truth (GT) illumination pool of each camera. The overall process is visualized in Fig. 8.
@@
While our camera-to-camera mapping is inspired by the C5 augmentation approach [afifi2021cross], it differs in the following ways. First, we remove C5’s restriction that limits sampling from the illumination pool to similar scenes with matching capturing settings (e.g., ISO, exposure time) and illumination CCT. Specifically, in C5, both the sampled scene image from the CIE XYZ space and the sampled illuminant from the target camera were required to have similar capturing settings and CCT. In contrast, our approach removes this constraint, eliminating the need to rely on capturing settings and allowing for greater diversity in augmentation. Additionally, instead of sampling from a fitted cubic polynomial based on the target camera’s illuminant samples, we use a fitted cubic polynomial based on the illuminant values from the source camera’s dataset (i.e., the camera from which the reference XYZ image was taken). The sampled illuminant is then transferred to the CIE XYZ space using the inverse of the source camera’s CM, followed by a transformation of these CIE XYZ illuminant values into the native raw RGB space of the target camera.
@@ C.3 Camera-specific Illumination Pool Generation
Next, we generate an illumination pool for each camera. While it is possible to use only the GT illuminations, we adopt the augmentation method proposed in [afifi2021cross] to enhance generality and diversity. This method involves fitting a cubic polynomial to the GT illuminations for each camera and then introducing random shifts to augment the illuminations. For further details, please refer to the supplementary material of [afifi2021cross]. On the right side of Fig. 8-(B), we show a plot of the illumination pool for a specific camera (Camera A). In this plot, the red points represent the GT illumination labels extracted from the dataset, while the blue points correspond to the augmented illuminations.
@@ Appendix E Experimental Setup
As mentioned in the main paper, the backbone $f$ uses the standard U-Net-like architecture from C5 [afifi2021cross]. However, unlike C5, we do not use additional images from the test camera, so no extra encoders are employed. Instead, we use a single Encoder-Decoder U-Net architecture. The encoder and decoder are connected via skip connections, with each consisting of four DoubleConv layers. In the encoder, each DoubleConv layer is followed by max pooling, while in the decoder, feature upsampling and skip connections are applied before each DoubleConv layer.
The batch size was set to 16, and training was conducted over 50 epochs with an initial learning rate of $5\times 10^{-4}$. A learning rate decay of 0.5 was applied at epoch 25. The Adam optimizer [kingma2014adam] was used for training.
For data augmentation, camera-to-camera mapping and imaginary camera augmentation are applied exclusively using the camera subsets from the training datasets, excluding the test dataset. For instance, when evaluating the Cube+ dataset, the augmented dataset used for model training is generated from images and CCMs from the Gehler-Shi [shi2000re], NUS-8 [cheng2014illuminant], and Intel-TAU [laakom2021intel] datasets.